a merchant has coffee worth 20 dollars a pound that she wishes to mix with with 60 pounds of coffee worth 90 dollars a pound to get a mixture that can be sold for eighty dollars a pound. How many pounds of twenty dollar offee should she use?
Answer by stanbon(75887)

a merchant has coffee worth 20 dollars a pound that she wishes to mix with with 60 pounds of coffee worth 90 dollars a pound to get a mixture that can be sold for eighty dollars a pound. How many pounds of twenty dollar offee should she use?
-----------------
Equation:
value + value = value
20x + 90*60 = 80(x+60) dollars
---
20x + 90*60 = 80x + 80*60
60x = 10*60
x = 10 lbs (amt of 20 dollar coffee needed in the mix)
--------------------------
Cheers,
Stan H.
================